OpenStudy (anonymous):

Good Morning, cans someone help me complete the square: 2x(x-3)=3x+2

OpenStudy (anonymous):

2x^2-6x=3x+2 2x^2-9x-2=0 x^2-9/2x-2=0 (x-9/4)^2-2+81/16=0 (x-9/4)^2+49/16=0

OpenStudy (anonymous):

okay. is it not necessary to square both sides?

OpenStudy (anonymous):

I dnt thnk u need to do t.

OpenStudy (anonymous):

my answer is \[ x= 9/4 \pm \sqrt{97} /4\]

OpenStudy (anonymous):

this is the format my instructor requires.

OpenStudy (anonymous):

(x-9/4)^2+49/16=0 (x-9/4)^2=-49/16 x=9/4+/-sqrt(-49)/4

OpenStudy (anonymous):

ok. thanks.

OpenStudy (anonymous):

u r welcum.